OpenStudy (anonymous):

PERFORM THE INDICATED OPERATIONS AND PLEASE SHOW WORK (5y - 7)^2

OpenStudy (diyadiya):

\[(a-b)^2=a^2+b^2-2ab\] \[(5y-7)^2=25y^2+49-70y\]

jhonyy9 (jhonyy9):

=(5y-7)(5y-7)=25y2 -70y +49